# Should We Still Pretrain Encoders with Masked Language Modeling?

### v2

@@ Abstract
Learning high-quality text representations is fundamental to a wide range of NLP tasks. While encoder pretraining has traditionally relied on Masked Language Modeling (MLM), recent evidence suggests that decoder models pretrained with Causal Language Modeling (CLM) can be effectively repurposed as encoders, often surpassing traditional encoders on text representation benchmarks. However, it remains unclear whether these gains reflect an inherent advantage of the CLM objective or arise from confounding factors such as model and data scale. In this paper, we address this question through a series of large-scale, carefully controlled pretraining ablations, training a total of 38 models ranging from 210 million to 1 billion parameters, and conducting over 15,000 fine-tuning and evaluation runs. We find that while training with MLM generally yields better performance across text representation tasks, CLM-trained models are more data-efficient and demonstrate improved fine-tuning stability. Building on these findings, we experimentally show that a biphasic training strategy that sequentially applies CLM and then MLM, achieves optimal performance under a fixed computational training budget. Moreover, we demonstrate that this strategy becomes more appealing when initializing from readily available pretrained CLM models, reducing the computational burden needed to train best-in-class encoder models. We release all project artifacts at https://hf.co/MLMvsCLM to foster further research.
@@ 1 Introduction
Figure 1: Experimental setup overview and key results on sequence classification (610M model size, 40% MLM ratio). The upper plot shows downstream performance as a function of the CLM–to-MLM step ratio during pretraining (configurations outperforming MLM-only are highlighted in yellow). The lower plot illustrates the effect of applying MLM CPT to models initially trained with either MLM or CLM (Base).
Learning meaningful representations is essential for a wide range of NLP tasks, such as sequence classification, named entity recognition, extractive question answering, and information retrieval. Traditionally, these representations are learned with encoders pretrained solely with Masked Language Modeling (MLM) (Devlin et al., 2019). More recently, however, decoder models pretrained with Causal Language Modeling (CLM) and later adapted with MLM have challenged the MLM-only paradigm (BehnamGhader et al., 2024), achieving state-of-the-art results on the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2023; Enevoldsen et al., 2025). Yet, these results have so far only been observed on models that are significantly larger than typical encoders and trained on substantially more data (Lee et al., 2024; Meng et al., 2024; Kim et al., 2024; Muennighoff et al., 2024). As a result, it remains unclear whether the success of this new training paradigm stems from the causal objective itself, or merely from increased scale.
In this paper, we conduct a controlled study to investigate the impact of training with MLM and CLM objectives when learning textual representations. We compare models with equal architecture and size, trained on equal amount of data, and evaluate them on an extensive range of text representation tasks. We start by comparing pretraining exclusively on either objective, and then investigate a two-stage pretraining alternative, where training starts with the CLM objective and is followed by MLM. Additionally, given the availability of existing pretrained models, we investigate a continued pretraining scenario, where models are initialized from checkpoints trained exclusively with MLM or CLM, and then trained with MLM for a fraction of the pretraining steps. In total, we train a total of 38 models ranging from 210 million to 1 billion parameters and conduct over 15,000 fine-tuning and evaluation runs, amounting to 110k MI250X GPU hours. Backed by these large-scale experiments, our results show that:
  • Although CLM training delivers strong performance on certain tasks and demonstrate good data efficiency and fine-tuning stability, MLM remains essential for achieving robust performance across all tasks, underscoring that bidirectional training remains indispensable (Section 3).
  • When pretraining a model from scratch, a two-stage protocol that first applies CLM followed by MLM offers a data-efficient way to build strong text representations by combining the strengths of both objectives (Section 4).
@@
To facilitate reproducibility and support future research, we release all pretrained model checkpoints, [^1] https://huggingface.co/MLMvsCLM along with the training [^2] https://github.com/Nicolas-BZRD/EuroBERT/tree/MLM_vs_CLM and evaluation code. [^3] https://github.com/hgissbkh/EncodEval/tree/MLM_vs_CLM
@@ 2 Experimental Setup
    CLM uses next-token prediction, where each token is predicted autoregressively using a causal attention mask. The training objective is to minimize the negative log-likelihood:
[…]
    where $x=(x_{1},x_{2},\ldots,x_{T})$ is the input sequence and $P(x_{t}\mid x_{<t})$ is the model’s predicted distribution over the vocabulary given the preceding tokens.
@@
    MLM, by contrast, randomly masks a subset of tokens and trains the model to reconstruct them using a bidirectional attention mask. The objective is:
[…]
    where $\mathcal{M}\subset\{1,\ldots,T\}$ denotes the indices of masked tokens, and $x_{\setminus\mathcal{M}}$ is the input sequence with masked tokens replaced by special placeholders. We experiment with masking ratios of 20%, 30%, 40%, and 50%.
@@
    Finally, the two-stage CLM+MLM approach sequentially applies CLM pretraining followed by MLM.
@@
Pretraining is performed with a per-device batch size of 12 samples across 192 GPUs, yielding an effective batch size of 2,373,120 tokens. [^5] Inputs consist of variable-length sequences ranging from 12 tokens up to the maximum of 2,048. We employ a Warmup-Stable-Decay (WSD) learning rate schedule: a 2,000-step warmup phase, followed by 38,000 steps with a constant learning rate of 5e-4, [^6] The optimal learning rate was selected from the range 1e-4 to 2e-3 based on experiments using 10% of the training data. ending with a 2,000-step linear decay phase, for a total of 42,000 training steps.
@@
To reflect two common real-world scenarios, we define two distinct pretraining setups, depending on whether the goal is to train a model from scratch or to continue training from an existing checkpoint:
@@
    Pretraining From Scratch (PFS): Models are trained from random initialization for a fixed number of steps using one of 3 objectives: CLM, MLM, or sequential CLM+MLM. For CLM and MLM, a standard WSD learning rate scheduler is applied. For CLM+MLM models, training is first performed with the CLM objective, and then resumed with MLM from CLM checkpoints that have not undergone learning rate decay.
@@
    Continued PreTraining (CPT): Models are initialized from existing checkpoints pretrained with either CLM or MLM, and training is resumed using the MLM objective. In contrast to the PFS setup, the pretrained models used for CPT have already undergone learning rate decay during their initial training phase, reflecting real-world constraints and continued training practices. Another key difference is that CPT starts from checkpoints where the loss has already converged, whereas in PFS, the objective switch typically occurs while gradient norms are still large and learning is active.
@@
We evaluate all models across a broad range of text representation tasks, focusing on 4 key categories commonly used in real-world applications. For Sequence Classification (SC), we use SST-2 (Socher et al., 2013), MNLI (Williams et al., 2018), and QQP (Wang et al., 2017). For Token Classification (TC), we evaluate on the English subsets of CoNLL (Tjong Kim Sang & De Meulder, 2003), OntoNotes (Hovy et al., 2006), and UNER (Mayhew et al., 2024). Question Answering (QA) is assessed using SQuAD (Rajpurkar et al., 2016), SQuAD-v2 (Rajpurkar et al., 2018), and ReCoRD (Wang et al., 2019). For Information Retrieval (IR), we use MS MARCO (Bajaj et al., 2016), NQ (Kwiatkowski et al., 2019), and the English subset of MLDR (Chen et al., 2024) for long-context evaluation. [^7] Further details are provided in Appendix B.
@@
To ensure a fair comparison, all models are fine-tuned using a consistent protocol. Each model is trained for up to 1,000 steps or one full epoch, whichever comes first, using a batch size of 32. To account for differences in model architecture and task requirements, we perform a grid search over 6 learning rates (1e-5, 2e-5, 5e-5, 1e-4, 2e-4, and 5e-4) for each model-dataset pair, with 10% warmup followed by linear decay. The learning rate yielding the best validation performance is selected. Fine-tuning employs bidirectional attention mask with task-specific loss functions: for SC, we use cross-entropy on mean-pooled token embeddings; TC and QA are trained using token-level cross-entropy; and for IR, we rely on the InfoNCE loss (Oord et al., 2018) with in-batch negatives, using mean pooling. To account for the fine-tuning instability commonly observed in BERT-style models for representation learning (Devlin et al., 2019; Lee et al., 2020; Dodge et al., 2020; Zhang et al., 2020), the entire procedure is repeated across 5 random seeds. Fine-tuning is conducted on the in-domain training set, except for NQ and MLDR, for which training is performed on MS MARCO.
Evaluation protocol.
@@
All pretraining and fine-tuning experiments are carried out on MI250X GPUs provided by the Adastra supercomputer, using the AMD-optimized EuroBERT codebase (Boizard et al., 2025).
@@
Drawing reliable conclusions about model design choices typically requires scaling up and repeating experiments, as statistically sound results often require a large number of runs and sufficient training to reach minimal convergence (Hoffmann et al., 2022). We therefore carefully design our experiments to ensure robust and well-supported findings. The default pretraining setup uses 100B tokens, which corresponds to 5 times the optimal data budget proposed by Hoffmann et al. (2022) for decoder models in the 1B-parameter range. Pretraining runs are executed on 24 nodes with 8 GPUs each (192 GPUs total), while fine-tuning is carried out on single GPUs. The total compute budget amounts to 110k GPU hours, broken down as follows:
  • Base pretraining includes all 3 model sizes (210M, 610M, and 1B), trained on both CLM and MLM objectives. For MLM, 4 masking ratios (20%, 30%, 40%, and 50%) are explored, resulting in a total of 15 models, each trained for 42,000 steps (100B tokens), and accounting for 81k GPU hours.
@@ 3 Pretraining with CLM or MLM
Figure 2: MLM vs. CLM downstream performance, averaged across tasks and reported for all model sizes. For MLM, results correspond to a 40% masking ratio.
Figure 3: Task-wise downstream performance across different masking ratios for all model sizes.
Figure 4: Downstream performance as a function of pretraining steps for CLM and MLM objectives. Results are reported for 610M models, with a 40% masking ratio for MLM.
Figure 5: Impact of the fine-tuning learning rate on MLM- vs. CLM-pretrained models. Error bars indicate the standard deviation of metric scores across all seeds and learning rates between 1e-5 and 1e-4. Results are shown for 610M-parameter models, with a 40% masking ratio for MLM.
This section presents preliminary results comparing the MLM and CLM pretraining objectives in terms of downstream performance, data efficiency, and training and fine-tuning stability.
@@
As expected, bidirectional attention during pretraining tends to enhance representation quality on downstream tasks (Figure 2). The performance gap between models trained with MLM and those trained with CLM is especially noticeable on SC and QA tasks, and remains consistent across model sizes. The largest discrepancy is observed on QA, which appears particularly sensitive to the absence of bidirectional attention during pretraining. Interestingly, some task-specific trends emerge: for instance, the MLM–to-CLM gap widens with increasing model size on SC, but narrows on IR.
@@
The masking ratio is a key design choice when pretraining models with MLM. As shown in Figure 3, there is no single best ratio, as it depends on both model size and downstream task, making MLM pretraining a delicate balance. Larger models tend to benefit from higher masking ratios, consistent with prior findings from Wettig et al. (2023). Across tasks, IR datasets consistently prefer higher masking ratios regardless of model size. In contrast, for token-level tasks such as TC and QA, smaller models perform better with lower masking ratios. For larger models (610M and 1B), the performance curves exhibit a U-shape, indicating improved performance at both low and high masking ratios. [^8] Additional results in Appendix C further highlight the influence of masking ratio, showing substantial variation even among datasets within the same task category. To reduce computational overhead, most subsequent experiments report results on 610M models trained with a 40% masking ratio for MLM, which provides a strong overall compromise across tasks.
@@
Interestingly, although CLM-pretrained models tend to underperform on SC, QA, and IR tasks, they achieve competitive results on TC, and even outperform MLM models at the 610M size (Figure 2). This suggests that causal pretraining can produce strong token-level representations, highlighting its potential despite the absence of bidirectional context.

### v3

@@ Abstract
Learning high-quality text representations is fundamental to a wide range of NLP tasks. While encoder pretraining has traditionally relied on Masked Language Modeling (MLM), recent evidence suggests that decoder models pretrained with Causal Language Modeling (CLM) can be effectively repurposed as encoders, often surpassing traditional encoders on text representation benchmarks. However, it remains unclear whether these gains reflect an inherent advantage of the CLM approach or arise from confounding factors such as model and data scale. In this paper, we address this question through a series of large-scale, carefully controlled pretraining ablations, training a total of 38 models ranging from 210 million to 1 billion parameters, and conducting over 15,000 fine-tuning and evaluation runs. We find that while training with MLM generally yields better performance across text representation tasks, CLM-trained models are more data-efficient and demonstrate improved fine-tuning stability. Building on these findings, we experimentally show that a biphasic training strategy that sequentially applies CLM and then MLM, achieves optimal performance under a fixed computational training budget. Moreover, we demonstrate that this strategy becomes more appealing when initializing from readily available pretrained CLM models, reducing the computational burden needed to train best-in-class encoder models. We release all project artifacts at https://hf.co/MLMvsCLM to foster further research.
@@ 1 Introduction
Learning meaningful representations is essential for a wide range of natural language processing (NLP) tasks, such as sequence classification, named entity recognition, extractive question answering, and information retrieval. Traditionally, these representations are learned with encoders pretrained solely with Masked Language Modeling (MLM) (Devlin et al., 2019). More recently, however, decoder models pretrained with Causal Language Modeling (CLM) and later adapted with MLM have challenged the MLM-only paradigm (BehnamGhader et al., 2024), achieving state-of-the-art results on the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2023; Enevoldsen et al., 2025). Yet, these results have so far only been observed on models that are significantly larger than typical encoders and trained on substantially more data (Lee et al., 2024; Meng et al., 2024; Kim et al., 2024; Muennighoff et al., 2024). As a result, it remains unclear whether the success of this training paradigm stems from CLM, or merely from increased scale.
In this paper, we present a controlled study examining the impact of training with MLM and CLM on learning textual representations. We compare models with equal architecture and size, trained on equal amount of data, and evaluate them on an extensive range of text representation tasks. We start by comparing pretraining exclusively with either approach, and then investigate a two-stage pretraining alternative, where CLM and MLM are applied sequentially. Additionally, leveraging existing pretrained models, we investigate a continued pretraining setup, where training starts from models initially trained with either MLM or CLM and is then extended with MLM for a portion of the pretraining steps. In total, we train 38 models ranging from 210 million to 1 billion parameters and conduct over 15,000 fine-tuning and evaluation runs to ensure fair design comparisons, amounting to 110k MI250X GPU hours. These large-scale experiments show that:
  • Although CLM training delivers strong performance on certain tasks and demonstrates good data efficiency and fine-tuning stability, MLM remains essential for achieving robust performance across all tasks, underscoring that bidirectional training remains essential (section 3).
  • When pretraining a model from scratch, a two-stage protocol that first applies CLM followed by MLM offers a data-efficient way to build strong text representations, effectively combining the strengths of both paradigms (section 4).
@@
To facilitate reproducibility and support future research, we release all pretrained model checkpoints, [^1] https://hf.co/MLMvsCLM together with the corresponding training [^2] https://github.com/Nicolas-BZRD/EuroBERT/tree/MLM_vs_CLM and evaluation codebases. [^3] https://github.com/artefactory/EncodEval/tree/MLM_vs_CLM
@@ 2 Experimental Setup
Figure 1: Experimental setup overview and key results on sequence classification (610M model size, 40% MLM ratio). The upper plot shows SC downstream performance as a function of the CLM–to-MLM step ratio during pretraining. The lower plot illustrates the effect of applying MLM CPT to models initially trained with either MLM or CLM (Base).
@@
    CLM employs next-token prediction, where each token is generated autoregressively using a causal attention mask. Given an input sequence $\mathbf{x}=(x_{1},x_{2},\ldots,x_{T})$, the training objective is to minimize the negative log-likelihood:
    $\mathcal{L}_{\text{CLM}}(\mathbf{x})=-\sum_{t=1}^{T}\log p_{\theta_{\rightarrow}}(x_{t}\mid x_{1},\ldots,x_{t-1}),$ (1)
    where $p_{\theta_{\rightarrow}}(\cdot\mid\cdot)$ denotes the model’s predicted distribution over the vocabulary conditioned on the preceding tokens under the causal attention mask.
@@
    MLM randomly masks a subset of tokens and trains the model to reconstruct them using a bidirectional attention mask. The sequence-level training objective is given by:
    $\mathcal{L}_{\text{MLM}}(\mathbf{x})=-\sum_{i\in\mathcal{M}}\log p_{\theta_{\leftrightarrow}}\!\left(x_{i}\mid\mathbf{x}_{\mathcal{M}}\right),$ (2)
    where $p_{\theta_{\leftrightarrow}}(\cdot\mid\cdot)$ denotes the model’s predicted distribution under a bidirectional attention mask, $\mathcal{M}\subset\{1,\ldots,T\}$ specifies the indices of the masked tokens, and $\mathbf{x}_{\mathcal{M}}$ is the input sequence with the corresponding tokens replaced by special placeholders. [^5] For example, if $T=5$ and $\mathcal{M}=\{1,3\}$, then $\mathbf{x}_{\mathcal{M}}=(\text{[mask]},x_{2},\text{[mask]},x_{4},x_{5})$. In practice, tokens are masked uniformly and independently, each with probability $p_{\text{mask}}$. In our experiments, we consider $p_{\text{mask}}\in\{20\%,30\%,40\%,50\%\}$.
@@
    Finally, CLM+MLM sequentially combines the former approaches in a two-stage setup, where CLM pretraining is performed first, followed by MLM.
@@
Pretraining is performed with a per-device batch size of 12 samples across 192 GPUs, yielding an effective batch size of 2,373,120 tokens. [^6] Inputs consist of variable-length sequences ranging from 12 tokens up to the maximum of 2,048. We employ a Warmup-Stable-Decay (WSD) learning rate schedule: a 2,000-step warmup phase, followed by 38,000 steps with a constant learning rate of $5\times 10^{-4}$, [^7] The optimal learning rate was selected from the range $1\times 10^{-4}$ to $2\times 10^{-3}$ based on experiments using 10% of the training data. ending with a 2,000-step linear decay phase, for a total of 42,000 training steps.
@@
We consider two pretraining setups reflecting common real-world scenarios: training a model from scratch and continuing training from an existing model.
@@
    Pretraining From Scratch (PFS). Models are trained from random initialization for a fixed number of steps using one of 3 approaches: CLM, MLM, or sequential CLM+MLM. For CLM and MLM, a standard WSD learning rate scheduler is applied. For CLM+MLM models, training is first performed with CLM, and then resumed with MLM from CLM checkpoints that have not undergone learning rate decay.
@@
    Continued PreTraining (CPT). Training is initialized from models pretrained using either CLM or MLM and is then continued with MLM. In contrast to the PFS setup, the pretrained models used for CPT have already undergone learning rate decay during their initial training phase, reflecting real-world constraints and continued training practices. Another key difference is that CPT starts from checkpoints where the loss has already converged, whereas in PFS, the paradigm switch typically occurs while gradient norms are still large and learning is active.
@@
We evaluate all models across a broad range of text representation tasks, focusing on 4 key categories commonly used in real-world applications. For Sequence Classification (SC), we use SST-2 (Socher et al., 2013), MNLI (Williams et al., 2018), and QQP (Wang et al., 2017). For Token Classification (TC), we evaluate on the English subsets of CoNLL (Tjong Kim Sang and De Meulder, 2003), OntoNotes (Hovy et al., 2006), and UNER (Mayhew et al., 2024). Question Answering (QA) is assessed using SQuAD (Rajpurkar et al., 2016), SQuAD-v2 (Rajpurkar et al., 2018), and ReCoRD (Wang et al., 2019). For Information Retrieval (IR), we use MS MARCO (Bajaj et al., 2016), NQ (Kwiatkowski et al., 2019), and the English subset of MLDR (Chen et al., 2024) for long-context evaluation. [^8] Further details are provided in Appendix B.
@@
To ensure a fair comparison, all models are fine-tuned using a consistent protocol. Each model is trained for up to 1,000 steps or one full epoch, whichever comes first, using a batch size of 32. To account for differences in model architecture and task requirements, we perform a grid search over 6 learning rates ($1\times 10^{-5}$, $2\times 10^{-5}$, $5\times 10^{-5}$, $1\times 10^{-4}$, $2\times 10^{-4}$, and $5\times 10^{-4}$) for each model-dataset pair, with 10% warmup followed by linear decay. The learning rate yielding the best validation performance is selected. Fine-tuning employs bidirectional attention mask with task-specific loss functions: for SC, we use cross-entropy on mean-pooled token embeddings; TC and QA are trained using token-level cross-entropy; and for IR, we rely on the InfoNCE loss (Oord et al., 2018) with in-batch negatives, using mean pooling. To account for the fine-tuning instability commonly observed in BERT-style models for representation learning (Devlin et al., 2019; Lee et al., 2020; Dodge et al., 2020; Zhang et al., 2020), the entire procedure is repeated across 5 random seeds. Fine-tuning is conducted on the in-domain training set, except for NQ and MLDR, for which training is performed on MS MARCO.
Evaluation metrics.
@@
All pretraining and fine-tuning experiments are carried out on MI250X GPUs provided by the Adastra supercomputer, using the AMD-optimized EuroBERT codebase (Boizard et al., 2025). Pretraining runs are executed on 24 nodes with 8 GPUs each (192 GPUs total), while fine-tuning is carried out on single GPUs.
@@
Drawing reliable conclusions about model design choices typically requires scaling up and repeating experiments, as statistically sound results often require a large number of runs and sufficient training to reach minimal convergence (Hoffmann et al., 2022). We therefore carefully design our experiments to ensure robust and well-supported findings. The default pretraining setup uses 100B tokens, which corresponds to 5 times the optimal data budget proposed by Hoffmann et al. (2022) for decoder models in the 1B-parameter range. The total compute budget amounts to 110k GPU hours, broken down as follows:
  • Base pretraining includes all 3 model sizes (210M, 610M, and 1B), trained with both CLM and MLM. For MLM, 4 masking ratios (20%, 30%, 40%, and 50%) are explored, resulting in a total of 15 models, each trained for 42,000 steps (100B tokens), and accounting for 81k GPU hours.
@@ 3 Pretraining with CLM or MLM
Figure 2: Downstream performance of MLM and CLM pretraining, averaged across tasks and reported for all model sizes. For MLM, results are given for a 40% masking ratio.
Figure 3: MLM downstream performance across different masking ratios for all model sizes.
Figure 4: Downstream performance with respect to pretraining steps for CLM and MLM. Results are reported for 610M models, with a 40% masking ratio for MLM.
Figure 5: Impact of the fine-tuning learning rate on MLM- vs. CLM-pretrained models. Error bars indicate the standard deviation of metric scores across all seeds and learning rates between $10^{-5}$ and $10^{-4}$. Results are shown for 610M-parameter models, with a 40% masking ratio for MLM.
This section presents preliminary results comparing MLM and CLM pretraining in terms of downstream performance, data efficiency, and training and fine-tuning stability.
@@
Bidirectional attention during pretraining tends to enhance representation quality on downstream tasks (Figure 2). The performance gap between models trained with MLM and those trained with CLM is especially noticeable on SC and QA tasks, and remains consistent across model sizes. The largest discrepancy is observed on QA, which appears particularly sensitive to the absence of bidirectional attention during pretraining. Interestingly, some task-specific trends emerge: for instance, the MLM–to-CLM gap widens with increasing model size on SC, but narrows on IR.
@@
The masking ratio is a key design choice when pretraining models with MLM. As shown in Figure 3, there is no single best ratio, as it depends on both model size and downstream task, making MLM pretraining a delicate balance. Larger models tend to benefit from higher masking ratios, consistent with prior findings from Wettig et al. (2023). Across tasks, IR datasets consistently prefer higher masking ratios regardless of model size. In contrast, for token-level tasks such as TC and QA, smaller models perform better with lower masking ratios. For larger models (610M and 1B), the performance curves exhibit a U-shape, indicating improved performance at both low and high masking ratios. [^9] Additional results in Appendix C further highlight the influence of masking ratio, showing substantial variation even among datasets within the same task category. [^10] In some configurations, we observe smaller models outperforming larger ones on average (e.g., the 210M model surpassing the 610M model on TC). Such behavior can arise from inherent pretraining variance (Grattafiori et al., 2024; Yang et al., 2025) and, in this case, results in no statistically significant differences. To reduce computational overhead, subsequent experiments report results on 610M models trained with a 40% masking ratio for MLM, which provides a strong overall compromise across tasks.
@@
Interestingly, although CLM-pretrained models typically underperform on SC and QA tasks, they achieve solid results on IR, with the gap to MLM models narrowing as model size increases. On TC, they consistently match MLM models and even outperform them by a substantial margin at the 610M scale (Figure 2). [^11] Intuitively, this behavior may stem from the nature of token-level tasks, which rely less on bidirectionally informed context to identify named entities or word types than SC, IR, or QA, which depend more heavily on broader contextual understanding. This hypothesis could be validated through further research. This suggests that causal pretraining can produce strong token-level representations, highlighting its potential despite the absence of bidirectional context.